# Prediction of the CP Phase $\delta_{CP}$ in the Neutrino Oscillation and an Axion-less Solution to the Strong CP Problem

### v2

@@
Abstract A model of the axion-less solution to the strong CP problem has been recently constructed based on the $\bf T^{2}/\mathbb{Z}_{3}$ orbifold in the six dimensional space-time. We extend, in this paper, the model to include the lepton sector with three right-handed neutrinos. A crucial point in the extended model is that we have only one CP-violating phase and thus the CP phase in the neutrino oscillation can be predicted by the CP phase in the CKM matrix. The CP phase $\delta_{CP}$ in the PMNS matrix is predicted as $\delta_{CP}\simeq 195^{\circ}-200^{\circ}$ which can be tested in the near future experiments. We also discuss the effective mass $m_{\beta\beta}$ responsible for the neutrino-less double beta decay of nucleus, which is predicted as $m_{\beta\beta}\simeq 8-11$ meV. )
@@ 1 Introduction
Recently, Liang, Okabe and one of the authors (T.T.Y.) have proposed a simple model for an axion-less solution to the strong CP problem [1], motivated by a consistent three zeros texture of the quark mass matrix in [2]. In this paper, we extend the original LOY model to the lepton sector including three right-handed neutrinos $N_{i}~{}(i$=1-3) and consider the leptogenesis mechanism for generating the baryon-number asymmetry[3] in the Universe. A surprising result in this extension is that we have only one CP-violation phase $\phi$ and the cosmological CP-violating phase is directly related to the CP phase $\delta^{\rm CKM}_{CP}$ in the CKM matrix [4]. Furthermore, we also show that the CP phase $\delta_{CP}$ in the neutrino oscillation is predicted by using the CP phase in the CKM matrix, which will be tested in near future experiments [5].
@@ 2 Extension of the LOY model to the lepton sector
$M_{D}=\begin{pmatrix}0&a^{\prime}_{\nu}&0\\ a_{\nu}&b_{\nu}\ e^{-i\phi}&c^{\prime}_{\nu}\\ 0&c_{\nu}&d_{\nu}\end{pmatrix}\,,$ (3)
@@
$\tilde{M}_{D}=\begin{pmatrix}0&A^{\prime}&0\\ A&B\ e^{-i\phi}&C^{\prime}\\ 0&C&D\end{pmatrix}\,.$ (4)
@@
$M_{\nu}=\begin{pmatrix}A^{\prime 2}&A^{\prime}B\,e^{-i\phi}&A^{\prime}C\\ A^{\prime}B\ e^{-i\phi}&A^{2}+C^{\prime 2}+B^{2}\,e^{-2i\phi}&BC\,e^{-i\phi}+C^{\prime}D\\ A^{\prime}C&BC\,e^{-i\phi}+C^{\prime}D\ &D^{2}+C^{2}\end{pmatrix}\,.$ (5)
@@
$\sin^{2}\theta_{12}$ $0.307^{+0.012}_{-0.011}$ $0.28-0.33$
$\sin^{2}\theta_{23}$ $0.454^{+0.019}_{-0.016}$ $0.42-0.59$
$\sin^{2}\theta_{13}$ $0.02224^{+0.00056}_{-0.00057}$ $0.021-0.023$
$\Delta m_{21}^{2}$ $7.41^{+0.21}_{-0.20}\times 10^{-5}{\rm eV}^{2}$ $(6.8-8.0)\times 10^{-5}\,{\rm eV}^{2}$
$\Delta m_{31}^{2}$     $2.505^{+0.024}_{-0.026}\times 10^{-3}{\rm eV}^{2}$ $(2.42-2.59)\times 10^{-3}\,{\rm eV}^{2}$
Table 3: The best fit ±$\pm 1\,\sigma$ and $2\sigma$ range of neutrino parameters from NuFIT 5.3 for NH (with SK atmospheric data) [6].
In addition, one should take into account the constraints from the CP Dirac phase $\delta_{CP}$ of NuFIT 5.3 [6] and the cosmological bound of the total sum of neutrino masses [12, 13, 14]. The CP Dirac phase is given by at the $2\sigma$ level:
$\delta_{CP}=185^{\circ}-320^{\circ}\,,$ (6)
@@ 3 Predictions of low-energy neutrino observables
Parameters of the neutrino mass matrix in Eq. (5) are determined by using the neutrino oscillation data in Table 3 with constraints of Eqs. (6) and (7)  [^5] The constraint Eq. (7) plays a crucial role for the determination of the six model parameters in the mass matrix $\tilde{M}_{D}$. since the charged lepton mass matrix in Eq. (2) is fixed. (In practice, the charged lepton masses are reproduced within $2\%$ error-bar at the electroweak scale.) We obtain the allowed ranges of parameters by scanning parameters in the real space with $|D|\geq|A|,\,|A^{\prime}|,\,|B|,\,|C|,\,|C^{\prime}|$ in order to realize NH of neutrino masses, where the absolute value of $D$ is fixed to reproduce the experimental data of $\Delta m^{2}_{31}$. The obtained numerical values is listed in Table 4. It is noticed that the absolute values of $A,\,A^{\prime},\,B,\,C$ are the same order to reproduce the large mixing angle between the first and second families, $\theta_{12}$, and $\Delta m^{2}_{21}$. On the other hand, $|C^{\prime}|$ is smaller than other parameters because the large mixing angle between the second and third families, $\theta_{23}$, is mostly furnished in the charged lepton mass matrix. [^6] In general, both large mixing angles between the second and third families in charged lepton mass matrix and the neutrino mass matrix can reproduce the observed $\sin\theta_{23}$ by tuning some phases. However, it is impossible in our scheme since we have only one fixed phase $\phi$.
@@
Because of the precise determination of the model parameters in Table 4, we can predict the Dirac CP phase $\delta_{CP}$, the total sum of neutrino masses and the effective mass $m_{\beta\beta}$ in the neutrino-less double beta decays of nucleus. At first, we present our prediction of the Dirac CP phase $\delta_{CP}$ versus $\sin^{2}\theta_{23}$ in Fig.4. The predicted range is $\delta_{CP}\simeq 195^{\circ}$-$200^{\circ}$, which is inside of $2\sigma$ range of NuFIT 5.3 (NH with SK atmospheric data) but outside its $1\sigma$ range [6]. The Hyper-Kamiokande can detect the signals in $3\sigma$ level in the future [5]. On the other hand, $\sin^{2}\theta_{23}$ is not allowed in the full range of $2\sigma$ in NuFIT 5.3 [6], but $\sin^{2}\theta_{23}\simeq 0.42-0.53$. This result will be tested in the near future.
@@
Figure 3: The predicted $\delta_{CP}$ versus $\sin^{2}\theta_{23}$. The region between the horizontal red lines denotes $1\sigma$ allowed one and the horizontal blue dashed line denotes the lower bound of $2\sigma$ allowed one in NuFIT 5.3 (NH with SK atmospheric data) [6]. The region between blue vertical dashed lines denotes $2\sigma$ allowed one.
@@ 4 The baryon-number asymmetry in the present Universe
${\rm Im}[(Y_{D}^{\dagger}Y_{D})^{2}_{12}]=-\frac{1}{v^{4}}\,a_{\nu}^{2}\,b_{\nu}^{2}\,\sin 2\phi\,,$ (11)
which is negative because of $\phi=37^{\circ}-48^{\circ}$ as seen in Table 2 and hence the $\epsilon_{1}$ is positive. Now, we see that our model predicts the wrong sign of the baryon-number asymmetry if the $N_{1}$ is the lightest RHN. However, if the lightest RHN is the $N_{2}$, the model predict the correct sign of the baryon asymmetry. [^7] We do not discuss, in this paper, the value of the asymmetry, since it depends on the details of the early universe such as the reheating processes and temperature and the correct mass of the lightest RHN. Important is here that the CP-violating phase for creating the baryon asymmetry is exactly the same as the CP-violating phase $\phi$ at low energies. Thus, it is extremely interesting that the presence of the baryon asymmetry in the Universe predicts the CP violation observed in the K-meson decay.
@@ 5 Conclusions
A simple model to solve the strong CP problem without introducing the axion has been recently proposed [1]. We consider, in this paper, an extension of the model to include leptons and the right-handed neutrinos (RHNs) $N_{i}~{}(i$=1-3). The remarkable point in this extension is that we have only one CP-violating phase $\phi$ and hence all CP-violation parameters in high-energy and low-energy processes are interconnected for each others. However, the sign of the cosmological baryon asymmetry depends also on mass order of the RHNs and we can not predict the sign of the cosmological baryon asymmetry as long as we do not know the mass hierarchy, as shown in section 4. We can, however, emphasise that the presence of the baryon asymmetry predicts strongly the CP violation in low energy processes.
@@
$\delta_{CP}\simeq 195^{\circ}-200^{\circ}\,.$

### v3

@@
Abstract A model of the axion-less solution to the strong CP problem has been recently constructed based on the $\bf T^{2}/\mathbb{Z}_{3}$ orbifold in the six dimensional space-time. We extend, in this paper, the model to include the lepton sector with three right-handed neutrinos. A crucial point in the extended model is that we have only one CP-violating phase and thus the CP phase in the neutrino oscillation can be predicted by the CP phase in the CKM matrix. The CP phase $\delta_{CP}$ in the PMNS matrix is predicted as $\delta_{CP}\simeq 192^{\circ}-195^{\circ}$ which can be tested in the near future experiments. We also discuss the effective mass $m_{\beta\beta}$ responsible for the neutrino-less double beta decay of nucleus, which is predicted as $m_{\beta\beta}\simeq 8-11$ meV. We consider the leptogenesis and confirm that our model generates the right sign of the baryon asymmetry if the first family right-handed neutrino is the lightest among the heavy right-handed neutrinos. )
@@ 1 Introduction
Recently, Liang, Okabe and one of the authors (T.T.Y.) have proposed a simple model for an axion-less solution to the strong CP problem [1], motivated by a consistent three zeros texture of the quark mass matrix in [2]. In this paper, we extend the original LOY model to the lepton sector including three right-handed neutrinos $N_{i}~{}(i$=1-3) and consider the leptogenesis mechanism for generating the baryon-number asymmetry[3] in the Universe. A surprising result in this extension is that we have only one CP-violating phase $\phi$ and the cosmological CP-violating phase is directly related to the CP phase $\delta^{\rm CKM}_{CP}$ in the CKM matrix [4]. Furthermore, we also show that the CP phase $\delta_{CP}$ in the neutrino oscillation is predicted by using the CP phase in the CKM matrix, which will be tested in near future experiments [5].
@@ 2 Extension of the LOY model to the lepton sector
$M_{D}=\begin{pmatrix}0&a^{\prime}_{\nu}&0\\ a_{\nu}&b_{\nu}\ {e^{i\phi}}&c^{\prime}_{\nu}\\ 0&c_{\nu}&d_{\nu}\end{pmatrix}\,,$ (3)
@@
$\tilde{M}_{D}=\begin{pmatrix}0&A^{\prime}&0\\ A&B\ {e^{i\phi}}&C^{\prime}\\ 0&C&D\end{pmatrix}\,.$ (4)
@@
$M_{\nu}=\begin{pmatrix}A^{\prime 2}&A^{\prime}B\,{e^{i\phi}}&A^{\prime}C\\ A^{\prime}B\ {e^{i\phi}}&A^{2}+C^{\prime 2}+B^{2}\,{e^{2i\phi}}&BC\,{e^{i\phi}}+C^{\prime}D\\ A^{\prime}C&BC\,{e^{i\phi}}+C^{\prime}D\ &D^{2}+C^{2}\end{pmatrix}\,.$ (5)
@@
$\sin^{2}\theta_{12}$ $0.308^{+0.012}_{-0.011}$ $0.28-0.33$
$\sin^{2}\theta_{23}$ $0.470^{+0.017}_{-0.013}$ $0.42-0.59$
$\sin^{2}\theta_{13}$ $0.02215^{+0.00056}_{-0.00058}$ $0.021-0.023$
$\Delta m_{21}^{2}$ $7.49^{+0.19}_{-0.19}\times 10^{-5}{\rm eV}^{2}$ $(7.11-7.87)\times 10^{-5}\,{\rm eV}^{2}$
$\Delta m_{31}^{2}$     $2.513^{+0.021}_{-0.019}\times 10^{-3}{\rm eV}^{2}$ $(2.47-2.56)\times 10^{-3}\,{\rm eV}^{2}$
Table 3: The best fit ±$\pm 1\,\sigma$ and $2\sigma$ range of neutrino parameters from NuFIT 6.0 for NH (with SK atmospheric data) [6].
In addition, one should take into account the constraints from the CP Dirac phase $\delta_{CP}$ of NuFIT 6.0 [6] and the cosmological bound of the total sum of neutrino masses [12, 13, 14]. The CP Dirac phase is given by at the $1\sigma$ level:
[…]
@@
The constraint Eq. (7) plays a crucial role for the determination of the six model parameters in the mass matrix $\tilde{M}_{D}$.
@@ 3 Predictions of low-energy neutrino observables
Parameters of the neutrino mass matrix in Eq. (5) are determined by using the neutrino oscillation data in Table 3 with constraints of Eqs. (6) and (7) since the charged lepton mass matrix in Eq. (2) is fixed. (In practice, the charged lepton masses are reproduced within $2\%$ error-bar at the electroweak scale.) We obtain the allowed ranges of parameters by scanning parameters in the real space with $|D|\geq|A|,\,|A^{\prime}|,\,|B|,\,|C|,\,|C^{\prime}|$ in order to realize NH of neutrino masses. In practice, we scan the ratios $A/D$, $A^{\prime}/D$, $B/D$, $C/D$ and $C^{\prime}/D$ in the range ±$\pm[0,1]$ with equal weights, that is the random scan in the linear space. The absolute value of $D$ is fixed to reproduce the experimental data of $\Delta m^{2}_{31}$. The obtained numerical values is listed in Table 4. It is noticed that the absolute values of $A,\,A^{\prime},\,B,\,C$ are the same order to reproduce the large mixing angle between the first and second families, $\theta_{12}$, and $\Delta m^{2}_{21}$. On the other hand, $|C^{\prime}|$ is smaller than other parameters because the large mixing angle between the second and third families, $\theta_{23}$, is mostly furnished in the charged lepton mass matrix. [^6] In general, both large mixing angles between the second and third families in charged lepton mass matrix and the neutrino mass matrix can reproduce the observed $\sin\theta_{23}$ by tuning some phases. However, it is impossible in our scheme since we have only one fixed phase $\phi$.
@@
Because of the precise determination of the model parameters in Table 4, we can predict the Dirac CP phase $\delta_{CP}$, the total sum of neutrino masses and the effective mass $m_{\beta\beta}$ in the neutrino-less double beta decays of nucleus.
[…]
At first, we present our prediction of the Dirac CP phase $\delta_{CP}$ versus $\sin^{2}\theta_{23}$ in Fig.4. The predicted range is $\delta_{CP}\simeq 192^{\circ}$-$195^{\circ}$, which is inside of $1\sigma$ range of NuFIT 6.0 (NH with SK atmospheric data) [6]. The Hyper-Kamiokande may detect the signals in $3\sigma$ level in the future [5]. On the other hand, $\sin^{2}\theta_{23}$ is not allowed in the full range of $2\sigma$ in NuFIT 6.0 [6], but $\sin^{2}\theta_{23}\simeq 0.42-0.53$. This result will be tested in the near future.
@@ 4 The baryon-number asymmetry in the present Universe
${\rm Im}[(Y_{D}^{\dagger}Y_{D})^{2}_{12}]={\frac{1}{v^{4}}\,a_{\nu}^{2}\,b_{\nu}^{2}\,\sin 2\phi}\,,$ (11)
which is positive because of $\phi=37^{\circ}-48^{\circ}$ as seen in Table 2 and hence the $\epsilon_{1}$ is negative. Now, we see that our model predicts the right sign of the baryon-number asymmetry if the $N_{1}$ is the lightest RHN. [^7] We do not discuss, in this paper, the value of the asymmetry, since it depends on the details of the early universe such as the reheating processes and temperature and the correct mass of the lightest RHN. Important is here that the CP-violating phase for creating the baryon asymmetry is exactly the same as the CP-violating phase $\phi$ at low energies. Thus, it is extremely interesting that the presence of the baryon asymmetry in the Universe predicts the CP violation observed in the K-meson decay.
@@ 5 Conclusions
A simple model to solve the strong CP problem without introducing the axion has been recently proposed [1]. We consider, in this paper, an extension of the model to include leptons and the right-handed neutrinos (RHNs) $N_{i}~{}(i$=1-3). The remarkable point in this extension is that we have only one CP-violating phase $\phi$ and hence all CP-violation parameters in high-energy and low-energy processes are interconnected for each others. However, the sign of the cosmological baryon asymmetry depends also on mass order of the RHNs and we can not predict the sign of the cosmological baryon asymmetry as long as we do not know the mass hierarchy, as shown in section 4. We can, however, emphasis that the presence of the baryon asymmetry predicts strongly the CP violation in low energy processes.
@@
$\delta_{CP}\simeq 192^{\circ}-195^{\circ}\,.$